# On the arithmetic and geometry of spaces $L_{m+1,n}$

### v1

@@ 3.2. A new generic obstruction to the existence of spaces $L_{\lambda p,2}$
For every $t\in k$, we denote by $Z_{t}$ be the set of zeroes of $P-tQ$ (counted with multiplicity). The condition $\deg(Q)<\deg(P)<p$ ensures that $Z_{D}$, $Z_{\delta}$ and $Z_{t}$ are all finite sets, as they are sets of roots of non-zero polynomials.
We have that $D(t)=c_{0}\mathrm{Res}(P-tQ,P^{\prime}-tQ^{\prime})$, where $c_{0}\in k$ does not depend on $t$. and we can then write
$D(t)=c_{1}\prod_{x\in Z_{t}}(P^{\prime}-tQ^{\prime})(x),$
for $c_{1}\in k^{\times}$ not depending on $t$. We have that $\gcd(P,Q)=1$, which guarantees that $\prod_{x\in Z_{t}}Q(x)\neq 0$, and we can then write
$D(t)=c_{1}\frac{1}{\prod_{x\in Z_{t}}Q(x)}\prod_{x\in Z_{t}}\big{(}P^{\prime}(x)Q(x)-tQ^{\prime}(x)Q(x)\big{)}=c_{2}\frac{1}{\mathrm{Res}(P-tQ,Q)}\mathrm{Res}(P^{\prime}Q-tQ^{\prime}Q,P-tQ),$
where $c_{2}\in k^{\times}$ is independent of $t$.
We now set $N:=P^{\prime}Q-PQ^{\prime}$ and denote by $Z_{N}$ its set of zeroes (counted with multiplicity). By using this notation and properties of resultants, we transform the equation above into
$D(t)=c_{2}\frac{1}{\mathrm{Res}(P,Q)}\mathrm{Res}(P^{\prime}Q-PQ^{\prime},P-tQ)=c_{3}\prod_{x\in Z_{N}}(P-tQ)(x),$
where $c_{3}\in k^{\times}$ is independent of $t$. Then, we define the following sets: the set $Z_{N,Q}$ of common zeroes of $N$ and $Q$ (counted with multiplicities) and the set ${}^{c}Z_{N,Q}$ of zeroes of $N$ that are not zeroes of $Q$. We note that, if $x\in Z_{N,Q}$ then it is also a zero of $Q^{\prime}$: as a result, its multiplicity as a zero of $Q$ is at least two, and precisely one more than the multiplicity of $x$ as a zero of $N$. From this, it follows that $Z_{N}=Z_{N,Q}\sqcup{}^{c}Z_{N,Q}$. On the other hand, if $x\in{}^{c}Z_{N,Q}$ we can use the fact that $\left(\frac{P}{Q}\right)^{\prime}=\frac{N}{Q^{2}}$ to deduce that the multiplicity of $x$ as a zero of $N$ is the same as the multiplicity as a zero of $\left(\frac{P}{Q}\right)^{\prime}$. It follows from this last consideration that $Z_{\delta}={}^{c}Z_{N,Q}$. In particular, $Z_{\delta}\cap Z_{Q}=\emptyset$ and the function $F$ is well defined.
The above equation then gets rewritten as
$D(t)$ $=c_{3}\prod_{x\in Z_{N,Q}}(P-tQ)(x)\prod_{x\in{}^{c}Z_{N,Q}}(P-tQ)(x)$
$=c_{3}\prod_{x\in Z_{N,Q}}P(x)\prod_{x\in{}^{c}Z_{N,Q}}(P-tQ)(x)$
$=c_{4}\prod_{x\in{}^{c}Z_{N,Q}}(P-tQ)(x)$
$=c_{4}\prod_{x\in{}^{c}Z_{N,Q}}Q(x)\prod_{x\in{}^{c}Z_{N,Q}}\left(\frac{P(x)}{Q(x)}-t\right)$
$=c_{5}\prod_{x\in{}^{c}Z_{N,Q}}\left(\frac{P(x)}{Q(x)}-t\right),$
where $c_{4},c_{5}\in k^{\times}$ are independent of $t$. It then follows that the function $F$ is surjective onto $Z_{D}$. From the equation, it also follows that $|Z_{D}|=|{}^{c}Z_{N,Q}|$. Since $Z_{\delta}={}^{c}Z_{N,Q}$, then $Z_{\delta}$ and $Z_{D}$ have the same cardinality and $F$ is bijective. ∎
@@
Assume by contradiction that there is a space $L_{\lambda p,2}$ with $p>3\lambda$. We set $P:=aP_{[1:0]}-a^{p}P_{[0:1]}$ and $Q:=a(P_{[1:0]}-P_{[0:1]})$ and we remark that these satisfy the conditions of Lemma 3.10. As a result, the zeroes of $z\mapsto D(z)=\mathrm{Disc}(P-zQ)$ are all of the form $\frac{P(x)}{Q(x)}$ for $x$ a zero of $\left(\frac{P}{Q}\right)^{\prime}$. In particular, $z=0$ is a zero of $D$ of order at most $\deg(P^{\prime})=\lambda-1$, as it corresponds to a zero $x$ of $\left(\frac{P}{Q}\right)^{\prime}$ that also satisfies $P(x)=0$ (and therefore also $P^{\prime}(x)=0$).
Let us now give a lower bound to the order of 00 as a zero of $D$ and see that it is incompatible with the one above. To do this, we apply the function $D$ to a new variable $z=\frac{t+a^{p}}{t+a}$. In this way, we have
@@
By Lemma 3.8 (ii), $-a^{p}$ is a zero of order at least $\lambda+3$ of the polynomial function $t\mapsto R(t)$ and, since the expression of $z$ in $t$ is a linear fractional transformation, we have that 00 is a zero of the same order of the polynomial $z\mapsto D(z)$. This gives the desired contradiction and concludes the proof of the theorem. ∎
@@ 4. Conditions for the existence of spaces $L_{\lambda p^{n-1},n}$
In this section, we prove a generalization of Proposition 3.1 that applies to spaces $L_{\lambda p^{n-1},n}$ for any $n\geq 2$ and discuss some of its consequences. As anticipated in 3.2, our strategy makes a crucial use of Moore determinants. Definition and results about Moore determinants that we use in this section are recalled in Appendix A. For every $n$-tuple of the form $\underline{X}:=(X_{1},\dots,X_{n})$, we denote by $\Delta_{n}(\underline{X})$ the associated Moore determinant. Moreover, we denote by $\underline{\hat{X}_{i}}$ the $n-1$-uple obtained from $\underline{X}$ by removing $X_{i}$ and by $\Delta_{n-1}(\underline{\hat{X}_{i}})$ the associated Moore determinant.
We develop our results in the following setting: we let $Q_{1},\dots,Q_{n}\in k[X]$ be polynomials of degree $\lambda\geq 1$ and denote by $q_{i}$ the leading coefficient of $Q_{i}$ for $1\leq i\leq n$. We write $P:=\Delta_{n}(\underline{Q})$, $P_{i}:=(-1)^{i-1}\Delta_{n-1}(\underline{\hat{Q}_{i}})$ and $P_{\underline{\epsilon}}:=\sum_{i}\epsilon_{i}P_{i}$ for every $\underline{\epsilon}\in{\mathbb{F}}_{p}^{n}-\{\underline{0}\}$.
@@
The polynomial $P$ and the $n$-tuple $\underline{P}=(P_{1},\dots,P_{n})$ satisfy the relation
@@
If we assume that $\Delta_{n}(\underline{q})\neq 0$, then the $q_{i}^{\prime}s$ are ${\mathbb{F}}_{p}$-linearly independent. This entails that $\deg P=(1+p+p^{2}+...+p^{n-1})\lambda$ and $\deg P_{\underline{\epsilon}}=(1+p+p^{2}+...+p^{n-2})\lambda$, since their leading coefficients are $\Delta_{n}(\underline{q})$ and $\sum_{i}\epsilon_{i}\Delta_{n-1}(\underline{\hat{q}_{i}})$ respectively. The first is nonzero by assumption, the second by virtue of [10, Corollary 2.1]
@@
Assume that $\Delta_{n}(\underline{q})\neq 0$. Then, we have that $P_{\underline{\epsilon}}|P$ for every $\underline{\epsilon}\in{\mathbb{F}}_{p}^{n}-\{\underline{0}\}$.
@@
and for every subspace $V\subset W$ we denote by $S(V)$ the intersection $V\cap S(W)$. It is a system of representatives of $\operatorname{Proj}(V)$. A counting argument shows that $\prod_{Q\in V-\{\underline{0}\}}Q=(-1)^{\dim V}(\prod_{Q\in S(V)}Q)^{p-1}$, which combined with the above gives the identity $P_{\underline{\epsilon}}^{p-1}=\left(\prod_{Q\in S(\ker\varphi_{\epsilon})}Q\right)^{p-1}$. We have then that there exists $\mu\in{\mathbb{F}}_{p}^{\times}$ such that $P_{\underline{\epsilon}}=\mu\prod_{Q\in S(\ker\varphi_{\underline{\epsilon}})}Q$. By Equation (A.2), we have that $P=\Delta_{n}(\underline{Q})=\prod_{Q\in S(W)}Q$. Since $S(\ker(\varphi_{\underline{\epsilon}}))\subset S(W)$ it follows that $P_{\underline{\epsilon}}|P$. ∎
@@
4.1. The main theorems
We now have all the tools to prove the two main results of this section (Theorems 4.7 and 4.8). Let us first establish some general results on the relationship between the $Q_{i}$’s and some spaces of differential forms that we can build from them:
@@ 4.1. The main theorems
For every $M\in GL_{n}({\mathbb{F}}_{p})$ we denote by $(\underline{Q}M)_{1},\dots,(\underline{Q}M)_{n}$ the components of the vector $\underline{Q}M$ obtained by applying the matrix $M$ to $\underline{Q}$. We note that all the entries $(\underline{Q}M)_{i}$’s are polynomials of degree $\lambda$ with leading coefficients that are ${\mathbb{F}_{p}}$-independent. We associate with this $n$-tuple the differential forms $\omega_{i}^{M}:=(-1)^{i-1}\frac{\Delta_{n-1}\big{(}\widehat{(\underline{Q}M)_{i}}\big{)}}{\Delta_{n}(\underline{Q}M)}dX$ and the space $\Omega^{\prime}:=\leavevmode\nobreak\ \langle\omega_{1}^{M},\dots,\omega_{n}^{M}\rangle_{{\mathbb{F}_{p}}}$.
@@
    $\omega_{j}^{M}:=$ $\frac{(-1)^{j+1}\Delta_{n-1}(\widehat{(\underline{Q}M)}_{j})}{\Delta_{n}(\underline{Q}M)}dX=\frac{(-1)^{j+1}\sum_{i=1}^{n}M_{i,j}\Delta_{n-1}(\underline{\hat{Q}_{i}})}{\Delta_{n}(\underline{Q}M)}dX=\frac{\sum_{i=1}^{n}{(-1)^{i+j}M_{i,j}P_{i}}}{\Delta_{n}(\underline{Q}M)}dX$
@@
Let $\Omega=\langle\omega_{1},\dots,\omega_{n}\rangle_{{\mathbb{F}_{p}}}$ be as in definition 4.4. For every $1\leq t\leq n$, let $\Omega_{t}\subset\Omega$ be the ${\mathbb{F}_{p}}$-subspace of $\Omega$ generated by $\{\omega_{n-t+1},\dots,\omega_{n}\}$ and let ${\mathcal{Q}}_{n-t}=\langle Q_{1},\dots,Q_{n-t}\rangle_{{\mathbb{F}_{p}}}$. Then, the $t$-uple of polynomials $\big{(}P_{{\mathcal{Q}}_{n-t}}(Q_{n-t+1}),\dots,P_{{\mathcal{Q}}_{n-t}}(Q_{n})\big{)}$ gives rise to the basis $(-1)^{n-t}(\omega_{n-t+1},...,\omega_{n}\big{)}$ of $\Omega_{t}$.
@@
As a result, the $t$-tuple $(P_{{\mathcal{Q}}_{n-t}}(Q_{n-t+1}),\dots,P_{{\mathcal{Q}}_{n-t}}(Q_{n}))$ gives rise to the basis of $\Omega_{t}$ given by the $t$-uple $\big{(}(-1)^{n-t}\omega_{n-t+1},\dots,(-1)^{n-t}\omega_{n}\big{)}$. ∎
@@
It follows that $\omega_{n-1}$ is also a logarithmic differential form. Using the same argument, we can show that $\omega_{i}$ is logarithmic for every $i=1,\dots,n-1$ and then that every form in $\Omega$ is logarithmic, which entails that $\Omega$ is a space $L_{\lambda p^{n-1},n}$.
To see that $P=\Delta_{n}(\underline{Q})$ has simple roots, note that its degree is precisely $\lambda\frac{p^{n}-1}{p-1}$, and that by construction its zeroes are the elements of ${\mathcal{P}}(\Omega)$, which we know by Lemma 2.3 to be a set of cardinality $\lambda\frac{p^{n}-1}{p-1}$. The roots then need to be simple. ∎
@@
Let $\Omega$ be a space $L_{\lambda p^{n-1},n}$ with $n\geq 2$. Then, there exist polynomials $Q_{1},\cdots,Q_{n}\in\leavevmode\nobreak\ k[X]$ of degree $\lambda$ such that, writing $P:=\Delta_{n}(\underline{Q})$ and $P_{i}:=(-1)^{i-1}\Delta_{n-1}(\underline{\hat{Q}_{i}})$, we have that $P_{i}|P$ and $\Omega=\langle\omega_{1},\dots,\omega_{n}\rangle_{{\mathbb{F}}_{p}}$ with $\omega_{i}=\frac{P_{i}}{P}dX$.
@@
The case $n=2$ is provided by Proposition 3.1, so we can proceed by induction: we fix $n\geq 3$ and assume that we have the result of the theorem in dimension up to $n-1$.
Let $(\omega_{1},\dots,\omega_{n})$ be a basis of $\Omega$, and let $P(X)=\prod_{x\in{\mathcal{P}}(\Omega)}(X-x)$. Then we can find $P_{1},\dots,P_{n}\in k[X]$ such that $\omega_{i}=\frac{P_{i}}{P}dX$. By Lemma 2.3 the degree of $P$ is $\lambda(1+p+\dots+p^{n-1})$ and then $\deg(P_{i})=\lambda(1+p+\dots+p^{n-2})$. By [10, Proposition 4.1] specialized at the case $q=p$ (cf. also the remark in [8, p. 68]) we have that $\Delta_{n}(P_{1},\dots,P_{n})=\gamma P^{1+p+\dots+p^{n-2}}$ for some $\gamma\in k^{\times}$. By possibly multiplying $P$ and the $P_{i}$’s by $\mu\in k^{\times}$ satisfying $\mu^{p^{n-1}}\gamma=1$, we can assume that

### v2

@@ 3.2. A new generic obstruction to the existence of spaces $L_{\lambda p,2}$
Let us write $N=P^{\prime}Q-PQ^{\prime}$, so that $\left(\frac{P}{Q}\right)^{\prime}=\frac{N}{Q^{2}}$. Clearly $(P\wedge P^{\prime})(Q\wedge Q^{\prime})$ divides $N$ and local computations at the zeroes of $PQ$, relying on the fact that both $\deg{(P)}$ and $\deg{(Q)}$ are prime to $p$, show that the quotient $N_{0}:=\frac{N}{(P\wedge P^{\prime})(Q\wedge Q^{\prime})}$ has no common roots neither with $P$ nor with $Q$. We can then set
$A=a(P\wedge P^{\prime})N_{0}\;\;\mbox{and}\;\;B=\frac{Q}{Q\wedge Q^{\prime}}$
for a suitable $a\in k^{\times}$ that makes $A$ monic. By construction, $A\wedge Q=1$ and $A\wedge B=1$. Hence, a zero of $\left(\frac{P}{Q}\right)^{\prime}$ is either a zero of $P\wedge P^{\prime}$ or of $N_{0}$.
Since $p$ does not divide $\deg(P)-\deg(Q)$, the term of degree $\deg P+\deg Q-1$ in $N$ is nonzero and hence $\deg(N)=\deg P+\deg Q-1$. Moreover, since $p$ does not divide $\deg(Q)$, one has $\deg(Q\wedge Q^{\prime})=\deg Q-\mathrm{rdeg}Q$. It follows that
$\deg(A)$ $=\deg((P\wedge P^{\prime})N_{0})=\deg(N)-\deg(Q\wedge Q^{\prime})=\deg P+\deg Q-1-(\deg Q-\mathrm{rdeg}Q)$
$=\deg P+\mathrm{rdeg}Q-1.$
Let us now prove the second part of the statement. For $t\in k$, we denote by $Z_{t}$ the set of zeroes of the polynomial $P-tQ\in k[X]$, and by $Z_{N}$ the set of zeroes of the polynomial $N$. We then compute $D(t)$ using properties of resultants: we have
$D(t)$ $=$ $\mathrm{Res}(P-tQ,P^{\prime}-tQ^{\prime})$
$=$ $c_{1}\hat{\prod}_{x\in Z_{t}}(P^{\prime}-tQ^{\prime})(x)$
where $c_{1}$ is independent of $t$, as it just involves the leading term of PtQP-tQ and we have by assumption that $\deg(Q)<\deg(P)$. As a result, we can write
$D(t)=c_{1}\frac{\hat{\prod}_{x\in Z_{t}}\left(P^{\prime}(x)Q(x)-tQ^{\prime}(x)Q(x)\right)}{\hat{\prod}_{x\in Z_{t}}Q(x)}.$
Note that this is possible as $P\wedge Q=1$ and then for every $x\in Z_{t}$ we have $Q(x)\neq 0$. Now, up to multiplication by a constant that does not depend on $t$, the numerator of the fraction above is equal to $\mathrm{Res}(P^{\prime}Q-tQ^{\prime}Q,P-tQ)$, while the denominator is equal to $\mathrm{Res}(P-tQ,Q)$. We can then write
$D(t)$ $=$ $c_{2}\frac{\mathrm{Res}(P^{\prime}Q-tQ^{\prime}Q,P-tQ)}{\mathrm{Res}(P-tQ,Q)}$
$=$ $c_{2}\frac{\mathrm{Res}(P^{\prime}Q-PQ^{\prime},P-tQ)}{\mathrm{Res}(P,Q)}$
$=$ $c_{3}\hat{\prod_{x\in Z_{N}}}(P-tQ)(x)$
where the new constant $c_{3}$ results from dividing $c_{2}$ by $\mathrm{Res}(P,Q)$ and multiplying by a suitable power of the leading term of $N$. Using this, we have
$D(t)$ $=$ $c_{3}\hat{\prod_{\begin{subarray}{c}x\in Z_{N}\\ Q(x)=0\end{subarray}}}(P-tQ)(x)\hat{\prod_{\begin{subarray}{c}x\in Z_{N}\\ Q(x)\neq 0\end{subarray}}}(P-tQ)(x)$
$=$ $c_{3}\hat{\prod_{\begin{subarray}{c}x\in Z_{N}\\ Q(x)=0\end{subarray}}}P(x)\hat{\prod_{\begin{subarray}{c}x\in Z_{N}\\ Q(x)\neq 0\end{subarray}}}(P-tQ)(x)$
$=$ $c_{4}\hat{\prod_{\begin{subarray}{c}x\in Z_{N}\\ Q(x)\neq 0\end{subarray}}}(P-tQ)(x)$
$=$ $c_{4}\hat{\prod_{\begin{subarray}{c}x\in Z_{N}\\ Q(x)\neq 0\end{subarray}}}Q(x)\hat{\prod_{\begin{subarray}{c}x\in Z_{N}\\ Q(x)\neq 0\end{subarray}}}\left(\frac{P}{Q}-t\right)(x)$
$=$ $c_{5}\hat{\prod_{\begin{subarray}{c}x\in Z_{N}\\ Q(x)\neq 0\end{subarray}}}\left(\frac{P(x)}{Q(x)}-t\right)$
$=$ $c_{5}\hat{\prod_{x\in Z_{A}}}\left(\frac{P(x)}{Q(x)}-t\right)$
where the final equation is justified by the facts that $N=a^{-1}A(Q\wedge Q^{\prime})$ and that $A\wedge Q=1$. ∎
@@
Assume by contradiction that there is a space $L_{\lambda p,2}$ with $p>3\lambda$. We set $P:=aP_{[1:0]}-a^{p}P_{[0:1]}$ and $Q:=a(P_{[1:0]}-P_{[0:1]})$ and we remark that these satisfy the conditions of Lemma 3.12. As a result, the zeroes of $D(Z)=\mathrm{Disc}(P-ZQ)\in k[Z]$ are all of the form $\frac{P(x)}{Q(x)}$ for $x$ a zero of $\left(\frac{P}{Q}\right)^{\prime}$. In particular, $z=0$ is a zero of $D$ of order at most $\deg(P^{\prime})=\lambda-1$, as it corresponds to a zero $x$ of $\left(\frac{P}{Q}\right)^{\prime}$ that also satisfies $P(x)=0$ (and therefore also $P^{\prime}(x)=0$).
Let us now give a lower bound to the order of 0 as a zero of $D$ and see that it is incompatible with the one above. To do this, we evaluate the polynomial $D$ at an element $z=\frac{t+a^{p}}{t+a}$. In this way, we have
@@
By Lemma 3.10 (ii), $a^{p}$ is a zero of order at least $\lambda+3$ of the polynomial function $t\mapsto R(t)$ and, since the expression of $z$ in $t$ is a linear fractional transformation, we have that 0 is a zero of the same order of the polynomial $D(Z)$. This gives the desired contradiction and concludes the proof of the theorem. ∎
Remark 3.13.
It is natural to ask whether the obstruction of Theorem 3.9 extends to the spaces $L_{\lambda p^{n-1},n}$. However, any $2$-dimensional subspace of a space $L_{\lambda p^{n-1},n}$ would have to satisfy the inequality $p>3\lambda p^{n-2}$ in order for Theorem 3.9 to apply and this condition is meaningful only when $n=2$. Consequently, no higher-dimensional obstructions arise directly from Theorem 3.9, and a more refined approach is required. We establish preliminary results in this direction—most notably Theorem 4.31 that settles the case $\lambda=1$ for every nn—but it remains open whether a generic condition yielding obstructions to the existence of spaces $L_{\lambda p^{n-1},n}$ for fixed $\lambda,n$, and almost every $p$ exists when $n>2$.
@@ 4. Conditions for the existence of spaces $L_{\lambda p^{n-1},n}$
In this section, we prove a generalization of Proposition 3.1 that applies to spaces $L_{\lambda p^{n-1},n}$ for any $n\geq 2$ and discuss some of its consequences. As anticipated in 3.4, our strategy makes a crucial use of Moore determinants. Definition and results about Moore determinants that we use in this section are recalled in Appendix A. For every $n$-tuple of the form $\underline{X}:=(X_{1},\dots,X_{n})$, we denote by $\Delta_{n}(\underline{X})$ the associated Moore determinant. Moreover, we denote by $\underline{\hat{X}_{i}}$ the n1n-1-tuple obtained from $\underline{X}$ by removing $X_{i}$ and by $\Delta_{n-1}(\underline{\hat{X}_{i}})$ the associated Moore determinant.
@@
Let $Q_{1},\dots,Q_{n}\in k[X]$ be polynomials of common degree $\lambda\geq 1$. Let $P:=\Delta_{n}(\underline{Q})$ and for every $i\in\{1,\dots,n\}$ let $P_{i}:=(-1)^{i-1}\Delta_{n-1}(\underline{\hat{Q}_{i}})$. Then, we have
@@
Let $Q_{1},\dots,Q_{n}\in k[X]$ be polynomials of common degree $\lambda\geq 1$. Let $P:=\Delta_{n}(\underline{Q})$ and let us define for every $i\in\{1,\dots,n\}$ the polynomials $P_{i}:=(-1)^{i-1}\Delta_{n-1}(\underline{\hat{Q}_{i}})$ and for every $\underline{\epsilon}\in{\mathbb{F}}_{p}^{n}-\{\underline{0}\}$ the polynomials $P_{\underline{\epsilon}}:=\sum_{i}\epsilon_{i}P_{i}$. Denote by $q_{i}$ be the leading coefficient of $Q_{i}$ and assume that $\Delta_{n}(\underline{q})\neq 0$. Then, the following conditions are met
  1. (i)
    We have the equality $\deg P=(1+p+p^{2}+\dots+p^{n-1})\lambda$
  2. (ii)
    For every $\underline{\epsilon}\in{\mathbb{F}}_{p}^{n}-\{\underline{0}\}$, $\deg P_{\underline{\epsilon}}=(1+p+p^{2}+\dots+p^{n-2})\lambda$
  3. (iii)
    For every $\underline{\epsilon}\in{\mathbb{F}}_{p}^{n}-\{\underline{0}\}$ $P_{\underline{\epsilon}}|P$.
@@
Let us prove the three items separately:
  1. (i)
    The leading coefficient of $P$ is $\Delta_{n}(\underline{q})$, which is non-zero by hypothesis, hence computing the Moore determinant gives $\deg P=(1+p+p^{2}+\dots+p^{n-1})\lambda$.
  2. (ii)
    The leading coefficient of $P_{\underline{\epsilon}}$ is $\sum_{i}\epsilon_{i}\Delta_{n-1}(\underline{\hat{q}_{i}})$, which is non-zero by [FresnelMatignon23, Corollary 2.1], hence $\deg P_{\underline{\epsilon}}=\deg{P_{1}}=(1+p+p^{2}+\dots+p^{n-2})\lambda$.
  3. (iii)
@@
    and for every subspace $V\subset W$ we denote by $S(V)$ the intersection $V\cap S(W)$. It is a system of representatives of $\operatorname{Proj}(V)$. A counting argument shows that $\prod_{Q\in V-\{\underline{0}\}}Q=(-1)^{\dim V}(\prod_{Q\in S(V)}Q)^{p-1}$, which combined with the above gives the identity $P_{\underline{\epsilon}}^{p-1}=\left(\prod_{Q\in S(\ker\varphi_{\epsilon})}Q\right)^{p-1}$. We have then that there exists $\mu\in{\mathbb{F}}_{p}^{\times}$ such that $P_{\underline{\epsilon}}=\mu\prod_{Q\in S(\ker\varphi_{\underline{\epsilon}})}Q$. By Equation (A.2), we have that $P=\Delta_{n}(\underline{Q})=\prod_{Q\in S(W)}Q$. Since $S(\ker(\varphi_{\underline{\epsilon}}))\subset S(W)$ it follows that $P_{\underline{\epsilon}}|P$.
@@
4.1. Constructing spaces $L_{\lambda p^{n-1},n}$ from polynomials
We now have all the tools to prove the two general results of this section (Theorems 4.7 and 4.10). Let us first establish some general results on the relationship between the $Q_{i}$’s and some spaces of differential forms that we can build from them:
@@ 4.1. Constructing spaces $L_{\lambda p^{n-1},n}$ from polynomials
For every $M\in GL_{n}({\mathbb{F}}_{p})$ we denote by $(\underline{Q}M)_{1},\dots,(\underline{Q}M)_{n}$ the components of the vector $\underline{Q}M$ obtained by applying the matrix $M$ to $\underline{Q}$. We note that all the entries $(\underline{Q}M)_{i}$’s are polynomials of degree $\lambda$ with leading coefficients that are $\mathbb{F}_{p}$}-independent. We associate with this $n$-tuple the differential forms $\omega_{i}^{M}:=(-1)^{i-1}\frac{\Delta_{n-1}\big(\widehat{(\underline{Q}M)_{i}}\big)}{\Delta_{n}(\underline{Q}M)}dX$ and the space $\Omega^{\prime}:=\penalty 10000\ \langle\omega_{1}^{M},\dots,\omega_{n}^{M}\rangle_{{\mathbb{F}_{p}}}$.
@@
    $\omega_{j}^{M}:=$ $\frac{(-1)^{j-1}\Delta_{n-1}(\widehat{(\underline{Q}M)}_{j})}{\Delta_{n}(\underline{Q}M)}dX=\frac{(-1)^{j-1}\sum_{i=1}^{n}M_{i,j}\Delta_{n-1}(\underline{\hat{Q}_{i}})}{\Delta_{n}(\underline{Q}M)}dX=\frac{\sum_{i=1}^{n}{(-1)^{i+j}M_{i,j}P_{i}}}{\Delta_{n}(\underline{Q}M)}dX$
@@
Let $\Omega=\langle\omega_{1},\dots,\omega_{n}\rangle_{{\mathbb{F}_{p}}}$ be as in definition 4.4. For every $1\leq t\leq n$, let $\Omega_{t}\subset\Omega$ be the $\mathbb{F}_{p}$}-subspace of $\Omega$ generated by $\{\omega_{n-t+1},\dots,\omega_{n}\}$ and let ${\mathcal{Q}}_{n-t}=\langle Q_{1},\dots,Q_{n-t}\rangle_{{\mathbb{F}_{p}}}$. Then, the $t$-tuple of polynomials $\big(P_{{\mathcal{Q}}_{n-t}}(Q_{n-t+1}),\dots,P_{{\mathcal{Q}}_{n-t}}(Q_{n})\big)$ gives rise to the basis $(-1)^{n-t}(\omega_{n-t+1},\dots,\omega_{n}\big)$ of $\Omega_{t}$.
@@
As a result, the $t$-tuple $(P_{{\mathcal{Q}}_{n-t}}(Q_{n-t+1}),\dots,P_{{\mathcal{Q}}_{n-t}}(Q_{n}))$ gives rise to the basis of $\Omega_{t}$ given by the $t$-tuple $\big((-1)^{n-t}\omega_{n-t+1},\dots,(-1)^{n-t}\omega_{n}\big)$. ∎
@@
It follows that $\omega_{n-1}$ is also a logarithmic differential form. Using the same argument, we can show that $\omega_{i}$ is logarithmic for every $i=1,\dots,n-1$ and then that every form in $\Omega$ is logarithmic, which entails that $\Omega$ is a space $L_{\lambda p^{n-1},n}$. ∎
Remark 4.8.
As a consequence of the theorem, we can prove that $P=\Delta_{n}(\underline{Q})$ has simple roots, and hence that the $Q_{i}$’s are pairwise coprime and have simple roots, generalising Remark 3.3 to any $n$. To see that $P=\Delta_{n}(\underline{Q})$ has simple roots, we note that its degree is precisely $\lambda\frac{p^{n}-1}{p-1}$, and that by construction its zeroes are the elements of ${\mathcal{P}}(\Omega)$, which we know by Lemma 2.3 to be a set of cardinality $\lambda\frac{p^{n}-1}{p-1}$, as $\Omega$ is a space $L_{\lambda p^{n-1},n}$. The roots of $P$ then need to be simple.
In analogy with the case $n=2$, Theorem 4.7 provides a directly computable method to verify that a space constructed as in Definition 4.4 is a space $L_{\lambda p^{n-1},n}$. Namely, one can apply Proposition 2.15 (iii) and conclude that it suffices to solve the following system of polynomial equations:
(4.9) $\begin{cases}\operatorname{coeff}\left(\big(\frac{P}{P_{n}}\big)^{p-1},X^{p-1}\right)=1\\ \operatorname{coeff}\left(\big(\frac{P}{P_{n}}\big)^{p-1},X^{\mu p-1}\right)=0,\;\;2\leq\mu\leq\lambda p^{n-2}(p-1).\end{cases}$
Note that, in adapting the proposition to the case of an $n$-dimensional space, the number of poles under consideration becomes $\lambda p^{n-1}$, and the number of equations must be computed accordingly.
@@
Let $\Omega$ be a space $L_{\lambda p^{n-1},n}$ with $n\geq 2$. Then, there exist polynomials $Q_{1},\cdots,Q_{n}\in\penalty 10000\ k[X]$ of degree $\lambda$ such that, writing $P:=\Delta_{n}(\underline{Q})$ and $P_{i}:=(-1)^{i-1}\Delta_{n-1}(\underline{\hat{Q}_{i}})$, we have that $P_{i}|P$ and $\Omega=\langle\omega_{1},\dots,\omega_{n}\rangle_{{\mathbb{F}}_{p}}$ with $\omega_{i}=\frac{P_{i}}{P}dX$.
@@
The case $n=2$ is provided by Proposition 3.1, so we can proceed by induction: we fix $n\geq 3$ and assume that we have the result of the theorem in dimension up to n1n-1.
Let $(\omega_{1},\dots,\omega_{n})$ be a basis of $\Omega$, and let $P(X)=\prod_{x\in{\mathcal{P}}(\Omega)}(X-x)$. Then we can find $P_{1},\dots,P_{n}\in k[X]$ such that $\omega_{i}=\frac{P_{i}}{P}dX$. By Lemma 2.3 the degree of $P$ is $\lambda(1+p+\dots+p^{n-1})$ and then $\deg(P_{i})=\lambda(1+p+\dots+p^{n-2})$. By [FresnelMatignon23, Proposition 4.1] specialized at the case $q=p$ (cf. also the remark in [Pagot02, p. 68]) we have that $\Delta_{n}(P_{1},\dots,P_{n})=\gamma P^{1+p+\dots+p^{n-2}}$ for some $\gamma\in k^{\times}$. By possibly multiplying $P$ and the $P_{i}$’s by $\mu\in k^{\times}$ satisfying $\mu^{p^{n-1}}\gamma=1$, we can assume that